Shift report 20000823 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 23 2000 - 00:09:02 EDT

  • Next message: http account: "Shift report 20000823 00:00-08:00"

    ------------------------------------------------------------
    Date: Wed Aug 23 00:09:02 2000, Shift: 16:00-24:00
    Supervisor: FR, with: ZM
    Log book#4 pages 142 - 143
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was cogged at 20:30 with I_b=14.7e9 and I_y=8.8e9.
       We noticed that trigger 1 rate was lower than expected
      (by comparing with #11 and #12).
    
    2. Run 2403 running since 20:30 
       - MRS at 90deg., D5 on (350A, A polarity).
       - FS at 5deg., D1 on (625A, B polarity), D2 on (492A,
         B polarity).
    
    3. T1 is off since last shift.
       No HV frame to bias T3.  
       Mo other detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 23 2000 - 00:12:38 EDT